### v1

# Dynamical stability for dense patterns in discrete attractor neural networks

Abstract
Neural networks storing multiple discrete attractors are canonical models of biological memory. Previously, the dynamical stability of such networks could only be guaranteed under highly restrictive conditions. Here, we derive a theory of the local stability of discrete fixed points in a broad class of networks with graded neural activities and in the presence of noise. By directly analyzing the bulk and outliers of the Jacobian spectrum, we show that all fixed points are stable below a critical load that is distinct from the classical critical capacity and depends on the statistics of neural activities in the fixed points as well as the single-neuron activation function. Our analysis highlights the computational benefits of threshold-linear activation and sparse-like patterns.
Introduction.
Attractor neural networks are canonical models of biological memory: they store neural activity patterns as stable fixed points (i.e., attractors) in their connection weights, so that when started from a noisy or incomplete version of one of these memory patterns as an initial condition, their autonomous dynamics converge to the corresponding fixed point owing to its dynamical stability – thus performing ‘auto-associative’ memory recall [1, 2]. Therefore, the stability of memory patterns as fixed points is critical for the operation of attractor networks. However, previous approaches had limited success in studying fixed-point stability.
The ‘Hebbian’ approach guarantees fixed-point stability in attractor networks by constructing an energy (or Lyapunov) function that is minimized by the network dynamics [1]. However, this has only been possible in a few (albeit very successful) cases [3, 4, 5, 6, 7] after making specific assumptions about the statistics of memory patterns (typically assumed to be binary), single-neuron activation functions (saturating, or rectified-linear), and in particular the way memory patterns influence dynamics (through some form of a so-called ‘Hebbian’ learning rule) requiring normal connection weight matrices. Even when those assumptions were violated, stability was achieved by approximately following such an energy function [8, 9, 10]. Conversely, the ‘Gardner’ approach allows the analysis of the storage capacity of neural networks in terms of the number of fixed points that can be embedded in their dynamics without recourse to an energy function [11, 12], but remains entirely mute about the stability of the embedded fixed points. Finally, optimization-based numerical approaches have also been used to embed stable fixed points in neural networks without making limiting assumptions, but they did not lend themselves to theoretical insight [13].
In this letter, we extend the ‘Gardner’ approach to gain analytical insights about the stability of fixed points in a broad class of networks with graded neural activities and generic, non-saturating, rectified, power-law activation functions. In particular, rather than the oft-studied sparse limit, here we consider dense patterns for analytical tractability on dynamical stability, and also because the inherent noisiness of neural signaling can easily prevent firing rates from being exactly zero in practice. To supplement these analyses, we also consider sparse patterns and show that our results extend to those in numerical simulations, and in some cases even analytically. We demonstrate that there is a phase transition for stability in such networks: optimizing network connectivity to maintain memory patterns as fixed points with minimal weights renders either all or none of those fixed points stable, depending on pattern statistics and single-neuron properties. We thus characterize the conditions under which fixed-point stability emerges in a large class of network dynamics, providing design principles for biological systems performing auto-associative memory [14, 15].
Figure 1: Storing fixed-points. (a) Color code for combinations of I/O function exponent and smoothness (log-scale). The activation function for selected combinations is illustrated (insets, top). These combinations are used in (b)-(d). (b,c) The sum (b) and L2-norm (c) of the presynaptic weights of a neuron (i.e. a row of $\mathbf{W}$) as a function of memory load ($\alpha$) for different activation functions (colors as in (a)), and $\textrm{CV}=1$. The inset shows divergence of L2-norm as $\alpha\rightarrow 1$, as predicted by theory (note linear scale for the y-axis and extended range on the x-axis). (d) The fraction of patterns which are correctly stored (thick curves) or stable for recall (thin curves) as a function of memory load ($\alpha$) for different activation functions (colors as in a-c), and $\textrm{CV}=2$. (e) Probability density of log-normal distributed patterns at different values of CV (color coded). The inset shows normalized density on a log scale. (f-h) Same as b-d for different values of pattern CV (colors as in e) and $\sigma=1$, $n=1.5$ (f,g); or $\sigma=1$, $n=1$ (h). Solid vs. dotted lines in (b)-(d), (f)-(h) show numerical vs. theoretical results. Numerical simulations optimized the weights $\mathbf{W}$ according to Eq. 4, and also optimized the threshold $\theta$ using finite differentiation. Note that only numerical results are shown for stability (d) and (h) – see subsequent figures for corresponding theoretical results.
Figure 2: Fixed-point stability. (a) Example spectrum - overlaid eigenvalues of $P$ Jacobians $\mathbf{J_{\mu}}$ of a single problem (parameters indicated by text). Annotations of the theoretical values $\lambda_{\mathrm{bulk}}$ (blue line), $\lambda_{\mathrm{mem}}$ (orange line), and $\lambda_{\mathrm{ave}}$ (green line). Inset shows a near-linear relation between $r^{\mu}_{i}$ (x-axis) and $f_{0}(r^{\mu}_{i})$ (y-axis), a prerequisite for an outlier $\lambda_{\mathrm{mem}}$. (b) Same as (a), for six parameter sets; columns differ by the maximal theoretical value. (c-e) Probability densities at different CV values (color-coded) - the difference between the empirical and predicted value of $\lambda_{\mathrm{bulk}}$ (c); the difference between the empirical spectral abscissa $\lambda^{*}$ (the real part of the eigenvalue with the largest real part) and the predicted value of $\lambda_{\mathrm{ave}}$ when varying the threshold $\theta$ such that $\lambda_{\mathrm{ave}}$ becomes the spectral abscissa (d); the difference between the empirical and predicted value of $\lambda_{\mathrm{mem}}$ (e).
Network model for auto-associative memory.
We study a network model of $N$ neurons with voltage dynamics:
$\tau\,\dot{\mathbf{v}}=-\mathbf{v}+\mathbf{W}\,\mathbf{g}\!\left(\mathbf{v}-\boldsymbol{\theta}\right)$ (1)
where $v_{i}$ is the voltage of neuron $i$, $\tau$ is the neural time constant (assumed to be shared across neurons), $\mathbf{W}\in\mathbb{R}^{N\times N}$ defines recurrent connection weights (such that $W_{ij}$ is the strength of the connection from neuron $j$ to neuron $i$), $g_{i}\!\left(\mathbf{v}\right)=g\!\left(v_{i}\right):\mathbb{R}\rightarrow\mathbb{R}^{+}$ is the neural activation function (also shared across neurons) that maps the ‘voltage’ of a neuron, $v_{i}$, to its (positive) instantaneous firing rate, and $\theta_{i}=\theta\in\mathbb{R}$ is the ‘threshold’ (or negative bias) for neuron $i$ (also shared across neurons). Defining $\mathbf{v}=\mathbf{W}\,\mathbf{r}$, Eq. 1 has an equivalent form of rate dynamics [16]
$\tau\,\dot{\mathbf{r}}=-\mathbf{r}+\mathbf{g}\!\left(\mathbf{W}\,\mathbf{r}-\boldsymbol{\theta}\right)$ (2)
which is the form we will use in the following for mathematical convenience.
While our theoretical results hold for a wide range of activation functions $g\!\left(\cdot\right)$, with additional assumptions detailed below, we consider here the specific case of the soft-rectified power-law to be able to systematically study how stability depends on its parameters (Fig. 1a):
$g\!\left(v\right)=\left[\frac{\sigma}{\pi}\,\ln\!\left(1+e^{\frac{\pi}{\sigma}\,v}\right)\right]^{n}$ (3)
with exponent $n$ and smoothness $\sigma$. The particular form of Eq. 3 is motivated as an approximation to a hard-rectified power-law activation function acting on noisy voltages, where voltage noise is Gaussian with standard deviation $\sigma$: $g\!\left(v\right)=\langle\lfloor v+\sigma z\rfloor^{n}_{+}\rangle_{z\sim{\cal N}\left(0,1\right)}$. Note that due to the noisiness of the activation function (or the equivalent smoothness in Eq. 3), all patterns in the network are dense as firing rates are never exactly zero. We also present results for a noiseless rectified power-law activation function that gives rise to sparse patterns (Details).
In line with previous approaches [11, 12], we begin by formalising the computational task for an auto-associative memory as the following: given a set of $P$ memory patterns $\mathbf{r}^{\mu}$ for $\mu=1\ldots P$, find minimal-norm weights such that each memory pattern is a fixed-point of the network dynamics (Eq. 2):
$\mathbf{W}^{*}=\mathop{\textrm{argmin}}_{\mathbf{W}}\left\|\mathbf{W}\right\|_{\mathrm{F}}\text{ s.t. }\mathbf{r}^{\mu}=\mathbf{g}\left(\mathbf{W}\,\mathbf{r}^{\mu}\right)\;\forall\mu=1\ldots P$ (4)
We consider the additional constraint that there are no self-couplings, i.e. the diagonal elements of the connection weight matrix are zero, $\mathrm{diag}\left(\mathbf{W}\right)=\boldsymbol{0}$.
We deviate from previous approaches [1, 2, 5, 12] in two important ways. First, we not only require memory patterns to be merely fixed points of the dynamics (which will be guaranteed once Eq. 4 is solved), but we also study the local stability of these fixed points – this is critical for a well-functioning auto-associative memory if it is to perform memory recall by pattern completion (converging to a memory pattern when started from a state that is near but not identical to it) [1, 17]. Specifically, we analyse the Jacobian of the dynamics around each memory pattern
$\mathbf{J}_{\mu}=-\mathbf{I}+\mathbf{g^{\prime}}\!\left(\mathbf{g}^{-1}\!\left(\mathbf{r}^{\mu}\right)\right)\circ\mathbf{W}$ (5)
The fixed-point at $\mathbf{r}^{\mu}$ is stable if all eigenvalues of the Jacobian have negative real parts.
[…]
To summarise, we are interested in how the maximal number of patterns $P$ that can be stored as fixed points, and the fraction of these fixed points that are dynamically stable, depends on parameters $\sigma$, $n$, $\theta$, and CV (as well as $f$, the sparseness of patterns, Details). For obtaining this maximum, we optimize $\mathbf{W}$ per Eq. 4.
Mean-field theory for storage capacity.
A replica analysis of the solution for the optimization problem Eq. 4 follows the Gardner approach [11, 12, 18], without assuming specific connection weights as in the Hopfield model [2, 5]. In the Appendix, we derive a mean-field theory for the capacity to store graded, random memory patterns as fixed points of Eq. 1, without assumptions on the weights. The analysis applies to either dense patterns with a strictly monotonic activation function or to sparse patterns with a rectified monotonic activation function. It generalizes previous work in which a rectified-linear activation function and normalized connection weights were assumed [12]. As usual, our analysis assumes $P,N\to\infty$ with a fixed memory load $\alpha=P/N$ and uses the replica technique and assuming the replica symmetry ansatz.
The resulting theory predicts that the critical capacity $\alpha_{\mathrm{C}}$ depends only on pattern sparseness and is independent of pattern statistics (CV; Fig. 1e) and activation function details (smoothness $\sigma$ and exponent $n$; Eq. 3, Fig. 1a), thus extending results from the rectified-linear case (corresponding to $\sigma=0$ and $n=1$) to our broader class of activation functions ($\sigma\geq 0$ and $n>0$) [12]. Results for the sparse case are presented in Fig. 5 (Details). In the dense case, the critical capacity is $\alpha_{\mathrm{C}}=1$ (note its independence from any of the parameters), and theory predicts the following moments for each element of the connection weight matrix:
$N\,\left\langle W^{2}\right\rangle$ $=\frac{\alpha}{1-\alpha}\frac{\left\langle\delta g^{-1}\!\left(r\right)^{2}\right\rangle}{\left\langle\delta r^{2}\right\rangle}$ (6)
$N\,\left\langle W\right\rangle$ $=\frac{\theta+\left\langle g^{-1}\!\left(r\right)\right\rangle}{\langle r\rangle}$ (7)
For a subcritical load, $\alpha=P/N<\alpha_{\mathrm{C}}$, theory can be compared with numerical experiments by directly solving Eq. 4 using off-the-shelf optimizers [19]. The theory provides a good match for both the mean ($N\,\left\langle W\right\rangle$ from Eq. 7; Fig. 1b,f and Fig. 5c) and the L2-norm of a row of the connection weight matrix ($\sqrt{N\,\left\langle W^{2}\right\rangle}$ from Eq. 6; Fig. 1c,g, and Fig. 5d). As $\alpha$ approaches $\alpha_{\mathrm{C}}$, there is no longer a solution to Eq. 4 (Fig. 1d,h and Fig. 5e), and the L2-norm of weights diverges, as predicted by theory (Fig. 1c,g, inset). Importantly, investigating the emergent stability of patterns in the subcritical regime reveals a second phase transition: all patterns tend to be stable up to a smaller critical value $\alpha_{\mathrm{S}}$, above which all patterns tend to be unstable (Fig. 1d,h and Fig. 6e,f). In the following, we investigate this phenomenon more closely.
Theory of stability for dense patterns.
For dense patterns, $\alpha_{\mathrm{C}}=1$ and for any $\alpha<\alpha_{\mathrm{C}}$ the solution to Eq. 4 is given in closed-form (Details):
$\mathbf{W}^{*}=\mathbf{V}\,\mathbf{R^{\dagger}}=\mathbf{V}\,\left(\mathbf{R}^{\mathsf{T}}\,\mathbf{R}\right)^{-1}\,\mathbf{R}^{\mathsf{T}}$ (8)
where $\mathbf{R},\mathbf{V}\in\mathbb{\mathbb{R}}^{N\times P}$ are the stored memory patterns $R_{i\mu}=r_{i}^{\mu}$ and $V_{i\mu}=g^{-1}(R_{i\mu})+\theta$. A slightly more complicated expression is available when avoiding self-couplings, Eq. 18, but we find that we can neglect this difference and use Eq. 8 in the derivation of stability. This generalizes earlier results derived for storing binary patterns, in which case weights with a similar structure, defined through the patterns’ pseudo-inverse, were already proposed in the ‘Hebbian’ approach[20, 3], but this was limited to a linear activation function $\mathbf{V}=\mathbf{R}$, rendering $\mathbf{W}$ symmetric (and thus allowing the construction of an energy function).
To characterize the critical load for stability, $\alpha_{\mathrm{S}}$, we previously analyzed the eigenvalue spectrum of random matrices $\mathbf{X}\,\mathbf{Y}^{\dagger}$ when pairs of the corresponding entries of $\mathbf{X}$ and $\mathbf{Y}$ are (jointly) independent and identically distributed (i.i.d.), correlated Gaussian [21]. This provides an exact result for the support of the eigenvalue spectrum (Fig. 7), and expressions for the largest real eigenvalue $\lambda_{\mathrm{bulk}}$ (Eq. 20), which we use to derive the critical load for stability, Eq. 21 (Details). For the Jacobian (Eq. 5) this is only an approximation, since the entries of the Jacobian are neither i.i.d. (due to the $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r^{\mu}}\right)\right)$ term which couples different rows), nor Gaussian (since the first term is a deterministic function of the second term, rather than being a jointly distributed random variable). Nevertheless, our empirical results below suggest that it is an acceptable approximation at least of the bulk of the eigenvalue spectrum, with some notable outliers that we will analyze separately below. The resulting critical load for stability from this analysis, describing the bulk of the eigenvalue spectrum, is:
$\alpha_{\mathrm{S}}^{\mathrm{bulk}}=\frac{\max\left(0,c_{\mathrm{rr}}-c_{\mathrm{rf}}\right)^{2}}{c_{\mathrm{rr}}\,c_{\mathrm{ff}}-c_{\mathrm{rf}}^{2}+\left(c_{\mathrm{rr}}-c_{\mathrm{rf}}\right)^{2}}$ (9)
expressed in terms of three cross-correlation terms $c_{\mathrm{rf}}=\left\langle\delta r\,\delta f\!\left(r\right)\right\rangle$, $c_{\mathrm{ff}}=\langle\delta f\!\left(r\right)^{2}\rangle$, and $c_{\mathrm{rr}}=\langle\delta r^{2}\rangle$, where $f\!\left(r\right)=\langle g^{\prime}\!\left(g^{-1}\!\left(r^{\prime}\right)\right)\rangle_{r^{\prime}}\left(g^{-1}\!\left(r\right)+\theta\right)$, with the averaging with respect to an independent variable $r^{\prime}$ neglecting the statistical dependence between different rows (i.e., ignoring the pattern-specific part of the Jacobian).
We note that $\alpha_{\mathrm{S}}^{\mathrm{bulk}}$ depends on $\theta$ only through $c_{\mathrm{ff}}$ and is maximized when $c_{\mathrm{ff}}$ is minimized, which occurs at:
$\theta^{*}=-\langle g^{-1}\!\left(r\right)\rangle$ (10)
so by Eq. 6 the average weight becomes 00 for this choice. Furthermore, we note that as long as the smoothness $\sigma$ is not too high, the threshold is expected to be negative.
The foregoing stability analysis was based on the zero-crossing of the ‘rightmost point’ of the bulk of the eigenvalue spectrum, $\lambda_{\mathrm{bulk}}$ (Fig. 2a). However, we empirically find that while $\lambda_{\mathrm{bulk}}$ often predicts stability well (Fig. 2c and Fig. 2b, left column), there are also cases when it alone is an imperfect predictor of stability (Fig. 2b, middle and right columns). Thus, we analyze two specific outlier eigenvalues that might interfere with stability.
The first outlier eigenvalue is associated with the uniform vector as a ‘counterfactual’ eigenvector. If such an eigenvector existed, the corresponding eigenvalue would be the (average) row-sum of the Jacobian, $\bar{J}$. The expected value of this outlier eigenvalue is thus:
$\lambda_{\mathrm{ave}}=\left\langle\bar{J}\right\rangle=-1+\left\langle g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)\right\rangle\,N\,\left\langle W\right\rangle$ (11)
where, by taking the expectation, we have ignored (again) the memory pattern-dependence of $g^{\prime}\!\left(g^{-1}\!\left(\mathbf{r}^{\mu}\right)\right)$ and $N\,\langle W\rangle$ can be computed using replica theory (Eq. 7):
$\lambda_{\mathrm{ave}}=-1+\left\langle g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)\right\rangle\,\frac{\theta+\left\langle g^{-1}\!\left(r\right)\right\rangle}{\langle r\rangle}$ (12)
Even though the uniform vector is not actually an eigenvector of the Jacobian in general, we empirically find that eigenvalue(s) often still exist close to $\lambda_{\mathrm{ave}}$ (Fig. 2a), and in fact these can even determine the dynamical stability of the Jacobian (Fig. 2d and Fig. 2b, middle column). Finally, requiring $\lambda_{\mathrm{ave}}<0$, Eq. 12 yields a constraint on the threshold $\theta$:
$\theta<-\langle g^{-1}\!\left(r\right)\rangle+\langle r\rangle/\langle g^{\prime}\!\left(g^{-1}\!\left(r\right)\right)\rangle$ (13)
which is always satisfied using $\theta^{*}$ from Eq. 10.
The second Jacobian matrix outlier is associated with the memory pattern itself. If for some constant $c$, $\mathbf{r^{\mu}}+c\mathbf{1}$ is an eigenvector of the Jacobian $\mathbf{J_{\mu}}$, it must satisfy:
$\mathbf{f_{0}\left(r^{\mu}\right)}=\left(\lambda_{\mathrm{mem}}^{\mu}+1\right)\mathbf{r^{\mu}}+c\left(\lambda_{\mathrm{mem}}^{\mu}-\lambda_{\mathrm{ave}}\right)\mathbf{1}$ (14)
where $f_{0}\left(x\right)=g^{\prime}\!\left(g^{-1}\!\left(x\right)\right)\left(g^{-1}\!\left(x\right)+\theta\right)$, and the linear relation between $\mathbf{r^{\mu}}$ and $\mathbf{f_{0}\left(r^{\mu}\right)}$ is captured by (Details):
$\tau_{\mathrm{mem}}^{\mu}$ $=\mathrm{corrcoef}\left(\mathbf{r^{\mu}},\mathbf{f_{0}\left(r^{\mu}\right)}\right)\approx 1$ (15)
$\lambda_{\mathrm{mem}}^{\mu}$ $=-1+c_{rf0}^{\mu}/c_{rr}$ (16)
with a new cross-correlation term $c_{rf0}^{\mu}=\left\langle\delta r^{\mu}\delta f_{0}\left(r^{\mu}\right)\right\rangle$. Empirically, it is straightforward to identify the eigenvector with the highest cross-correlation to the memory pattern (associated eigenvalue marked in orange at Fig. 2a), systematically assess the validity of the estimator $\lambda_{\mathrm{mem}}$ (Fig. 2e) and find cases where they determine the dynamical stability of the Jacobian (Fig. 2b, right column).
We note an interesting symmetry: the bulk-related instability $\lambda_{\mathrm{bulk}}\geq 0$ when $c_{rf}\geq c_{rr}$, and the pattern-specific instability $\lambda_{\mathrm{mem}}^{\mu}\geq 0$ when $c_{rf0}^{\mu}\geq c_{rr}$. The former was derived by neglecting pattern-specific statistical dependency between Jacobian rows, while the latter is a direct result of this dependency.
Importantly, the two outliers $\lambda_{\mathrm{ave}},\lambda_{\mathrm{mem}}$ depend only on the pattern statistics and the activation function, but not on $\alpha$, i.e., on the number of stored patterns, so our estimate for the critical load for stability:
$\alpha_{\mathrm{S}}=\alpha_{\mathrm{S}}^{\mathrm{bulk}}\Theta\left(-\lambda_{\mathrm{mem}}\right)\Theta\left(-\lambda_{\mathrm{ave}}\right)\Theta\left(\tau_{\mathrm{mem}}-0.9\right)$ (17)
To show the qualitative and quantitative predictions of the theory more systematically, we numerically analyzed the critical load for stability for different combinations of pattern statistics CV, activation function smoothness $\sigma$, and exponent $n$. Fig. 3 present the results as phase diagrams in the $\sigma$ vs $n$ plane. Critical stability is zero in many areas on this plane, with the stable regime defined by the intersection of 3 conditions: (1) $\lambda_{\mathrm{bulk}}>0$ prevents stability at most of the sublinear regime; (2) $\lambda_{\mathrm{mem}}>0$ prevents stability at most of the supralinear regime; (3) $\tau_{\mathrm{mem}}<0.9$ at large values of $\sigma$, invalidating the assumption of Eq. 15; so instability is not captured by Eq. 16. We thus predict a surprising mechanism for instability for $n>1$, due to $\lambda_{\mathrm{mem}}>0$ (Fig. 3). Intuitively, in this regime, a perturbation in the direction of the memory pattern, or scaling it up by a factor, will be amplified due to a supralinear exponent and diverge.
Figure 3: Phase diagram of stability. (a-e) The predicted value of $\alpha_{\mathrm{S}}$ (color coded) for different combinations of I/O exponent (x-axis) and I/O smoothness (y-axis, log-scale). A plus sign marks the location of the maxima. Pattern variation level (CV) is indicated at each panel. Boundaries of the different regimes are marked: $\lambda_{\mathrm{bulk}}>0$ (blue line), $\lambda_{\mathrm{mem}}>0$ (orange line), $\tau_{\mathrm{mem}}<0.9$ (red line). (a) also overlays a grid of numerical results: yellow disk marks stability (with size-coded value), gray cross marks stability below detectable range (most crosses removed to avoid clutter).
Figure 4: Optimal parameters. (a-b) The optimal exponent (a) or smoothness (b), at different pattern variation levels (x-axis). (c) The critical load for stability using optimal parameters, at different pattern variation levels (x-axis). (d) Measures of the weights asymmetry (dark) and dynamics non-normality (light) at different pattern variation levels (x-axis). The asymmetry index is $\left\|\mathbf{W}_{\mathrm{asym}}\right\|_{\mathrm{F}}/\left(\left\|\mathbf{W}_{\mathrm{sym}}\right\|_{\mathrm{F}}+\left\|\mathbf{W}_{\mathrm{asym}}\right\|_{\mathrm{F}}\right)$ for the symmetric and anti-symmetric parts of the weights $\mathbf{W}$. The non-normality index is Henrici’s deviation from normality index [22, 23] for the Jacobian $\mathbf{J}$, $\sqrt{\left\|\mathbf{J}\right\|_{\mathrm{F}}^{2}-\sum_{i}\left|\lambda_{i}\right|^{2}}/\left\|\mathbf{J}\right\|_{\mathrm{F}}$. Presenting the median value across choices of exponent and smoothness.
Optimal parameter.
The overall comparison of theory and numerical experiments is presented in Fig. 8. While the match is not perfect, Fig. 3 explains the sources of the discrepancy. The points with numerical results close to zero but substantial prediction from theory, or vice versa, all come from the $n\approx 1$ region, where any imprecision in estimating the boundary $\lambda_{\mathrm{mem}}=0$ changes the prediction between an order-one value and 0.
Fig. 4a-b shows the optimal exponent and smoothness for different levels of pattern variation. Our analysis predicts substantial stability only in the near-linear region $n\approx 1$, and at a finite value of $\sigma$. The benefits of threshold-linear activation are also observable for sparse patterns (Fig. 6e), where it is optimal for $\textrm{CV}\geq 1$. This is consistent with previous literature on storage of graded memory patterns using the ‘Hebbian’ approach, which focused on this activation function [5, 12] and with the machine learning literature using similar activations [24], but contrasts with other approaches which advocates for the benefits of supralinear activation [25, 26].
Theory further predicts improvement of stability with pattern variation (Fig. 4c). Those high-variance memory patterns are sparse-like, with many near-zero values and few large activations, so the improved capacity we report for those echoes the benefit of sparseness in our results (Fig. 6e) and previous works [5, 4]. Such sparse-like patterns are also consistent with the known phenomenology of neurons in the brain [27, 28].
To test if the optimal weights used here may coincide with those assumed by the ‘Hebbian’ approach, we measured the weight matrix asymmetry and the deviation from normality of the dynamics around the fixed points (Fig. 4d). Both measures tend to increase with variation levels, indicating that a high critical load for stability is likely achieved for asymmetric weights, and leads to dynamics distinct from those of energy-based models.
The resulting optimal threshold is usually negative in our theory for dense patterns (Eq. 10), and also empirically for sparse patterns (Fig. 6c). A negative threshold is uncommon in the literature (a notable exception is the Willshaw model [29]) and contrasts with the prevailing view in the field [17], which holds that the construction of multiple distinct and stable activity patterns requires global inhibition and selective recurrent excitation.
Discussion.
Our analysis highlights the conditions for memory pattern stability in the auto-associative memory task. The resulting theory is rich and makes many testable predictions, most notably a stability phase transition at a number of patterns which is proportional to the number of neurons, and is below the capacity phase transition, which does not account for stability. It constrains the range of usable activation thresholds and predicts the optimal single-neuron activation function for different memory pattern statistics. Most notably, we establish that near-linear activation and finite noise levels are optimal for stable recall of dense memory patterns.
Those insights are relevant to neuroscience and cannot be derived from the non-biological, energy-based, auto-associative memory networks commonly used in the field. This letter opens the door for designing networks with biological-relevant dynamics capable of storing many patterns as stable fixed points, with simple optimization. Future research may establish if biological-like local learning can be used for such optimization.
Many areas of science study dynamical systems where fixed points and their dynamic stability may be of interest [30]. Theoretically, this question may be analyzed through the eigenvalue spectrum of the dynamics’ Jacobian at the fixed point. Full characterization of the spectrum is possible for many classes of random matrices [31, 32, 33, 34, 35, 36, 37] and exhibits a universality property [38]. Many random dynamical systems exhibit a stability phase transition when the largest real value in the eigenvalues spectrum crosses zero [30, 39, 36, 40, 41]. Those random matrices can have an exponential number of fixed points, with stability which depends on a global criterion [42, 35, 43]. In this letter, we go beyond random matrices (and low-rank perturbations thereof [44, 45]) to describe fixed-point stability in a system resulting from a learning or an optimization process. In such systems, the connection weight matrix (and subsequently the Jacobian) is defined in terms of the constraints it satisfies or the objective function it minimizes. The stability of fixed points in a learned network could be previously analyzed only for very simple cases, such as a scalar output [46]. Even for the heavily studied Hopfield model, the stability of patterns remains unclear, and memory patterns are not, in general, fixed points of the dynamics [2]. Our analysis of the eigenvalue spectrum of the Jacobian may be applicable to non-random systems in other fields, when the optimal connectivity is given by a pseudo-inverse rule. Admittedly, this analysis needs to be complemented with a problem-specific approach to characterize outliers.
Acknowledgements.
This work was supported by the Wellcome Trust (Investigator Award in Science 212262/Z/18/Z to M.L.), the Human Frontiers Science Programme (Research Grant RGP0044/2018 to M.L.), and the Blavatnik Cambridge Postdoctoral Fellowships (to U.C.).